OpenStudy (anonymous):

what is the gcd(235,50)

OpenStudy (anonymous):

235= 4(50)+35 so the gcd is 200

OpenStudy (anonymous):

mmm sry now 35= 4(8) + 3 so th gcd = 3 that's it